Find, showing all working, a recursive definition of the sequence with general term

tn = 6 (n + 1)!/3n, n >= 1


1
Expert's answer
2021-10-11T05:49:49-0400

Let us find a recursive definition of the sequence with general term

"t_n = \frac{6 (n + 1)!}{3^n}, n \ge 1."

It follows that

"t_{n+1} = \frac{6 (n + 2)!}{3^{n+1}}=\frac{6 (n + 1)!(n+2)}{3^{n}\cdot3}=\frac{6 (n + 1)!}{3^{n}}\frac{n+2}{3}=t_n\frac{n+2}{3}."

We conclude that the recursive definition of this sequence is

"t_{n+1}=\frac{1}3(n+2)t_n,\ t_1=4."